OpenStudy (anonymous):

The point P (pi/2 , 0) lies on the curve y = xcosx . Find the equation of the normal to the curve at the point P.

OpenStudy (isaiah.feynman):

You can't assume the line is perfectly perpendicular to the curve so we assume the following. Assume another line is TANGENT to the curve and passes through P(pi/2, 0). Find the slope of that tangent line by taking the derivative of y=xcosx, then substitute pi/2 for x in the derivative, the new value gotten is the SLOPE of the tangent line. The normal is perfectly perpendicular to the tangent line and we know that the product of the slopes of perpendicualr lines is -1 so use m1m2=-1 to find the slope of the normal line m1 is the slope of the tangent, so solve for m2.

OpenStudy (anonymous):

First of all, how do I differentiate xcosx?

OpenStudy (anonymous):

We use the multiplication rule \[ \frac{d}{dx}(x*Cosx)=1*Cos(x)+x*(-Sinx)= Cosx-x*Sinx\]

OpenStudy (anonymous):

So when P = (pi/2), dy/dx = cos pi/2 - pi/2 sin pi/2

OpenStudy (anonymous):

= -1.57?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Is it alright to leave it as -1.57 or does it have to be -pi/2?

OpenStudy (anonymous):

Well its easier to leave it as -pi/2 Cos(pi/2)=0 and sin(pi/2)=1 So \[ \cos( \frac{\pi}{2})-\frac{\pi}{2} \sin(\frac{\pi}{2})= 0-\frac{\pi}{2}(1)=-\frac{\pi}{2}\]

OpenStudy (anonymous):

So obviously the slope of the normal is \(\large \frac{ 2}{\pi}\) since \[ \frac{ 2}{\pi}*-\frac{ \pi}{2}=-1\]

OpenStudy (anonymous):

Why is it 2/pi again? I forgot

OpenStudy (isaiah.feynman):

pi/2*m2=-1, what is m2 @Lii

OpenStudy (anonymous):

I don't know lol

OpenStudy (anonymous):

-0.63?

OpenStudy (isaiah.feynman):

If we have the equation. \[-\frac{ \pi }{ 2 } \times m_{2} =-1\] How can you solve for m2?

OpenStudy (anonymous):

Make m2 the subject

OpenStudy (anonymous):

so 0.63 so therefore 2/pi

OpenStudy (isaiah.feynman):

Exactly! that answers your question.

OpenStudy (anonymous):

Ok so y-0 = 2/pi (x-pi/2)

OpenStudy (anonymous):

y = 2/pi x - 1

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

Thanks but as -pi/2 = -1.57, would substituting -1.57 still get the right answer?

OpenStudy (anonymous):

@Isaiah.Feynman so if I substituted -1.57 instead: y-0 = -1.57(x-pi/2) y-0 = -1.57x + 2.46 y = -1.57x + 2.46, is that right or wrong?

OpenStudy (isaiah.feynman):

I think it would be wrong.

OpenStudy (anonymous):

Oh ok, well thanks @Isaiah.Feynman and @BlackLabel